How many grams of NaHCO3 (sodium bicarbonate, or baking soda) would you need to make a 2 molar (2M) solution?

How many grams of C12H22O11 (sucrose) would you need to make a 0.5 molar (1M) solution?
1
Expert's answer
2019-10-01T08:06:55-0400


(i)

Let's say the solution volume is 1 L. Then:

"n(NaHCO_3) = c*V"


"n(NaHCO_3) = 2 M *1 L = 2mol"

"m(NaHCO_3) = n*M"

"m(NaHCO_3) = 2mol*84g\/mol = 168 g."

(ii)

Let's say the solution volume is 1 L. Then:

"n(C_{12}H_{22}O_{11}) = c * V"

If solution 0.5 molar, then "n(C_{12}H_{22}O_{11}) = 0.5 mol".

Then "m(C_{12}H_{22}O_{11}) = n*M = 0.5mol*342g\/mol = 171 g".

If solution is 1M, then "n(C_{12}H_{22}O_{11}) = 1 mol".

Then "m(C_{12}H_{22}O_{11}) = 342g."